OpenStudy (anonymous):

what value of c makes the expression a perfect square? x^2+x+c

OpenStudy (anonymous):

c=1/4

OpenStudy (anonymous):

howd u get that?

OpenStudy (anonymous):

\[(x + a)^2 = x^2 + 2ax + a^2 = x^2 + x + c\]so\[2a = 1 \Rightarrow a = \frac{1}{2} \Rightarrow c = a^2 = \frac{1}{4}.\]